There were 80 runners to start a race. In the first half of the race, 2/5 of them dropped out. In the second half of the race, 3/8 of the remaining runners dropped out. How many runners finished the race?
1
Expert's answer
2013-03-29T07:27:20-0400
In the first half of the race, 2/5 of all runners droppedout.
Therefore it remains
80*(1-2/5) = 80*3/5 = 48 runners.

In the second half of the race, 3/8 of the remainingrunners dropped out, so it remains
48*(1-3/8) = 48*5/8 = 30 runners.

Thus, only 30 runners finished the race